# Primal Methods for Variational Inequality Problems with Functional Constraints

### v1

@@ Abstract
Constrained variational inequality problems are recognized for their broad applications across various fields including machine learning and operations research. First-order methods have emerged as the standard approach for solving these problems due to their simplicity and scalability. However, they typically rely on projection or linear minimization oracles to navigate the feasible set, which becomes computationally expensive in practical scenarios featuring multiple functional constraints. Existing efforts to tackle such functional constrained variational inequality problems have centered on primal-dual algorithms grounded in the Lagrangian function. These algorithms along with their theoretical analysis often require the existence and prior knowledge of the optimal Lagrange multipliers. In this work, we propose a simple primal method, termed Constrained Gradient Method (CGM), for addressing functional constrained variational inequality problems, without necessitating any information on the optimal Lagrange multipliers. We establish a non-asymptotic convergence analysis of the algorithm for variational inequality problems with monotone operators under smooth constraints. Remarkably, our algorithms match the complexity of projection-based methods in terms of operator queries for both monotone and strongly monotone settings, while utilizing significantly cheaper oracles based on quadratic programming. Furthermore, we provide several numerical examples to evaluate the efficacy of our algorithms.
@@ 1 Introduction
Here, the constrained set $\mathcal{C}\subseteq\mathbb{R}^{d}$ is compact and convex, and the operator $F:\mathcal{C}\rightarrow\mathbb{R}^{d}$ is continuous, which ensures a nonempty and compact solution set [15]. In this work, we also assume the operator $F$ to be monotone, a standard assumption in the literature that encompasses important applications including convex minimization and convex-concave minimax optimization problems [3, 55, 56, 47].
There has been a growing interest in the development of first-order methods for solving monotone variational inequality problems, e.g., extragradient [32, 62, 47, 7] and optimistic gradient methods [54, 41, 40, 20]. These methods, noted for their simplicity and scalability, involve querying the operator $F$ at specific points and accessing a projection oracle or a linear minimization oracle [24, 16] for the feasible set $\mathcal{C}$. Such oracles are easy to compute for simple feasible sets such as Euclidean balls for the projection oracle or polytopes for the linear minimization oracle. However, for general constrained sets, computing the projection or linear minimization oracle requires solving constrained optimization problems over the feasible set $\mathcal{C}$. The latter remains challenging even with quadratic or linear objectives.
@@
Primal methods serve as an alternative approach to avoid such explicit dependence on the optimal Lagrange multipliers. They are also simpler to analyze and more straightforward to implement in practice. In fact, primal first-order methods have been extensively studied for functional constrained minimization problems. Prominent examples include Polyak’s switching gradient method [52, 33, 22], cutting-plane methods [29], the level bundle method [34], and the constrained gradient descent method [42] motivated from nonsmooth dynamical systems, to just name a few. However, primal methods have been much less explored for functional constrained variational inequality problems.
@@
We establish the global convergence analysis of CGM under two settings: $(i)$ when the operator $F$ is monotone, and $(ii)$ when $F$ is strongly-monotone. Notably, in both settings, CGM enjoys (nearly) the same complexity on querying the operator $F$ as the optimal complexity achieved by projection-based methods, even though the projection oracle is replaced by quadratic programming solvers. For example, when the operator $F(x)$ is monotone, we show that CGM achieves an $\epsilon$-solution with $\mathcal{O}(1/\epsilon^{2})$ queries to $F(x)$ and $\mathcal{O}(1/\epsilon^{2})$ calls to a quadratic programming solver. To the best of our knowledge, our algorithm is the first primal method that achieves the optimal complexity on queries to $F$ for functional constrained monotone variational inequality problems without dependence on the optimal Lagrange multipliers.
@@
We further illustrate that the quadratic programming at each iteration of CGM allows for efficient implementation and permits even closed-form solutions in special cases such as simplex constraints (Algorithm 4) or when there is only one (nonlinear) constraint function (Algorithm 2). This yields a direct implementation of our algorithm comparable to unconstrained methods. Empirically, we evaluate the convergence of CGM through several numerical experiments and demonstrate its effectiveness.
@@ 2.1 Variational Inequality Problems
In our paper, we assume that $\mathcal{C}$ is contained in an Euclidean ball with diameter $D$. This suggests that the feasible set $\mathcal{C}$ is convex and compact. By Corollary 2.2.5 in Facchinei and Pang [15], the solution set of the variational inequality problem (1) is nonempty and compact. We are interested in finding the following $\epsilon$-approximate solution of the constrained variational inequality problem (1).
@@
For some $\epsilon>0$, we call a point $\hat{x}\in\mathbb{R}^{d}$ an $\epsilon$-solution of the variational inequality problem (1), where the feasible set is of the form (2), if $F(x)^{\top}(\hat{x}-x)\leq\epsilon$, $\forall x\in\mathcal{C}$ and $g_{i}(\hat{x})\leq\epsilon$, $\forall i\in[m]$.
The above definition is often referred to as the weak $\epsilon$-solution of a variational inequality problem and is commonly adopted in the related literature [47, 48, 28, 5]. A strong $\epsilon$-solution, $\hat{x}\in\mathcal{C}$, is characterized by $F(\hat{x})^{\top}(\hat{x}-x)\leq\epsilon$, $\forall x\in\mathcal{C}$. When $\epsilon=0$ and $F$ is monotone and continuous, the two solution sets are equivalent [39, 30]. When $\epsilon>0$ and $F$ is monotone, every strong $\epsilon$-solution is also a weak $\epsilon$-solution.
@@ 3 The Constrained Gradient Method (CGM)
A standard algorithm in the literature on constrained monotone variational inequality problems is the projected gradient method, which requires access to a projection oracle and performs the following updates at each iteration $t=0,1,\ldots,T-1$:
@@
In Algorithm 1, we allow for an inexact solution of the quadratic program, where the stopping criteria can be satisfied using various methods. More details on how this quadratic program can be solved are discussed in Section 5. The auxiliary constraint $g_{m+1}(x)=\lVert x\rVert^{2}-D^{2}$ is used to ensure that both the iterate $x_{t}$ and the projected velocity $v_{t}$ are bounded. It is added specifically to simplify the theoretical analysis and will not be used in actual implementations.
@@ 4.1 The Monotone Setting
The above theorem implies that for CGM to achieve an $\epsilon$-solution, the number of queries to the operator $F(x)$ is at most $\mathcal{O}(1/\epsilon^{2})$, and the number of calls to a quadratic programming solver is at most $\mathcal{O}(1/\epsilon^{2})$. Notably, the complexity of querying the operator $F$ matches that of projection-based methods [49] or Frank-Wolfe type methods [61], known to be optimal even for simple constraint sets according to the lower complexity bound (Theorem 3.2.1 in Nesterov [49]).
In stark contrast, the convergence rates of previous primal-dual methods exhibit explicit dependence on the optimal Lagrange multipliers. For instance, the ADMM-based interior point method [66, 8] achieves strong $\epsilon$-solutions with the same $\mathcal{O}(1/\epsilon^{2})$ queries to $F$ and requires $\mathcal{O}(1/\epsilon^{2})$ calls to a program that solves a sequence of strongly-convex minimization problems. However, their results require the strong assumption that either $F$ is strictly-monotone or one of the $g_{i}(x)$ is strictly-convex. The constraint extrapolation (ConEx) method [5, 6] achieves an $\epsilon$-solution as in Definition 1 with $\mathcal{O}(1/\epsilon^{2})$ calls to the operator $F$. However, they also require the existence and boundedness of the optimal Lagrange multipliers, whose magnitude explicitly appears in the convergence upper bounds for both optimality and feasibility gaps. To the best of our knowledge, CGM is the first primal method that achieves the optimal complexity on queries to $F$ for functional constrained monotone variational inequality problems without dependence on the optimal Lagrange multipliers.
@@ 4.2 The Strongly-Monotone Setting
When $F(x)$ is $\mu$-strongly-monotone, i.e., $(F(x_{t})-F(x))^{\top}(x_{t}-x)\geq\mu\lVert x_{t}-x\rVert^{2}$, we obtain that $\forall\,x\in\mathcal{C}$,
@@
$(t+2)^{c_{\gamma}}g_{i}(x_{t+1})$ $\leq\left(1-\frac{c_{\gamma}}{t+1}\right)(t+2)^{c_{\gamma}}g_{i}(x_{t})+\frac{c_{M}(t+2)^{c_{\gamma}}}{(t+1)^{2}}$
$=\left(1-\frac{c_{\gamma}}{t+1}\right)\left(1+\frac{1}{t+1}\right)^{c_{\gamma}}(t+1)^{c_{\gamma}}g_{i}(x_{t})+\frac{c_{M}}{(t+1)^{2-c_{\gamma}}}\left(1+\frac{1}{t+1}\right)^{c_{\gamma}}$
$\leq\left(1-\frac{c_{\gamma}}{t+1}\right)\left(1+\frac{c_{\gamma}}{t+1}\right)(t+1)^{c_{\gamma}}g_{i}(x_{t})+\frac{c_{M}}{(t+1)^{2-c_{\gamma}}}\left(1+\frac{c_{\gamma}}{t+1}\right)$
$\leq(t+1)^{c_{\gamma}}g_{i}(x_{t})+\frac{2\,c_{M}}{(t+1)^{2-c_{\gamma}}},$ (9)
@@
$(\kappa(t+1)+2)^{c_{\gamma}}g_{i}(x_{\kappa(t+1)+1})$
$\quad\leq(\kappa(t+1)+2)^{c_{\gamma}}\max\left\{\frac{ML_{g}}{\kappa(t+1)+1},\frac{c_{M}}{(\kappa(t+1)+1)^{2}}\right\}$
$\quad=\left(1+\frac{1}{\kappa(t+1)+1}\right)^{c_{\gamma}}\max\left\{\frac{ML_{g}}{(\kappa(t+1)+1)^{1-c_{\gamma}}},\frac{c_{M}}{(\kappa(t+1)+1)^{2-c_{\gamma}}}\right\}$
$\quad\leq 2\max\left\{\frac{ML_{g}}{(\kappa(t+1)+1)^{1-c_{\gamma}}},\frac{c_{M}}{(\kappa(t+1)+1)^{2-c_{\gamma}}}\right\}$
$\quad\leq 2\max\{ML_{g},c_{M}\}.$
@@
In the strongly-monotone setting, the convergence rate of CGM in terms of the optimality gap is $\mathcal{O}(1/T)$, which matches the rate of projection-based methods [49]. However, it is worth noting that the rate in terms of the constraint violation approaches $\mathcal{O}(1/T)$ only as $\gamma$ grows large enough. We leave the task of closing this gap to future work. As a comparison, the primal-dual method in Yang et al. [66] achieves $\mathcal{O}(1/\sqrt{T})$ rate measured by the distance $\lVert\bar{x}_{T}-x^{*}\rVert$ and does not provide explicit feasibility guarantees. If the operator is additionally Lipschitz, the same $\mathcal{O}(1/\sqrt{T})$ rate is attained for a strong approximate solution.

### v2

@@ Abstract
Variational inequality problems are recognized for their broad applications across various fields including machine learning and operations research. First-order methods have emerged as the standard approach for solving these problems due to their simplicity and scalability. However, they typically rely on projection or linear minimization oracles to navigate the feasible set, which becomes computationally expensive in practical scenarios featuring multiple functional constraints. Existing efforts to tackle such functional constrained variational inequality problems have centered on primal-dual algorithms grounded in the Lagrangian function. These algorithms along with their theoretical analysis often require the existence and prior knowledge of the optimal Lagrange multipliers. In this work, we propose a simple primal method, termed Constrained Gradient Method (CGM), for addressing functional constrained variational inequality problems, without requiring any information on the optimal Lagrange multipliers. We establish a non-asymptotic convergence analysis of the algorithm for Minty variational inequality problems with monotone operators under smooth constraints. Remarkably, our algorithms match the complexity of projection-based methods in terms of operator queries for both monotone and strongly monotone settings, while using significantly cheaper oracles based on quadratic programming. Furthermore, we provide several numerical examples to evaluate the efficacy of our algorithms.
@@ 1 Introduction
Here, the constrained set $\mathcal{C}\subseteq\mathbb{R}^{d}$ is compact and convex, and the operator $F:\mathcal{C}\rightarrow\mathbb{R}^{d}$ is continuous, which ensures a nonempty and compact solution set [23]. The definition above is also known as the Stampacchia variational inequality problem, and the solution $x^{*}$ is referred to as a strong solution to the variational inequality problem corresponding to $F$ and $\mathcal{C}$. Our aim is to find a weak solution $x^{*}\in\mathcal{C}$ to the variational inequality problem, where it solves the Minty variational inequality problem such that $F(x)^{\top}(x^{*}-x)\leq 0,\forall\,x\in\mathcal{C}$. In this work, we also assume the operator $F$ to be monotone, a standard assumption in the literature that encompasses important applications including convex minimization and convex-concave minimax optimization problems [5, 78, 79, 66]. More applications on monotone variational inequality problems can be found in Thekumparampil et al. [87] and Chapter 1.4 of Facchinei and Pang [23]. When $F$ is continuous and monotone, strong and weak solution sets are equivalent [58, 44, 19], and weak solutions are commonly adopted in the related literature for monotone variational inequality problems [66, 67, 41].
There has been a growing interest in the development of first-order methods for solving monotone variational inequality problems, e.g., the projected gradient method [23], the proximal point method [79], the extragradient method [46, 88, 66, 13], the optimistic gradient method [77, 60, 59, 30], the dual extrapolation method [67, 69], and the projected reflected gradient method [55]. These methods, noted for their simplicity and scalability, involve querying the operator $F$ at specific points and accessing a projection oracle or a linear minimization oracle [36, 24] for the feasible set $\mathcal{C}$. Such oracles are easy to compute for simple feasible sets such as Euclidean balls for the projection oracle or polytopes for the linear minimization oracle. However, for general constrained sets, computing the projection or linear minimization oracle requires solving constrained optimization problems over the feasible set $\mathcal{C}$. The latter remains challenging even with quadratic or linear objectives.
@@
Primal methods serve as an alternative approach to avoid such information on the optimal Lagrange multipliers. They are also simpler to analyze and more straightforward to implement in practice. In fact, primal first-order methods have been extensively studied for functional constrained minimization problems. Prominent examples include Polyak’s switching gradient method [76, 49, 34], cutting-plane methods [43], the level bundle method [51], and the constrained gradient descent method [61] motivated from nonsmooth dynamical systems, to just name a few. However, primal methods have been much less explored for functional constrained variational inequality problems.
@@
We establish the global convergence analysis of CGM under two settings: $(i)$ when the operator $F$ is monotone, and $(ii)$ when $F$ is strongly-monotone. Notably, in both settings, CGM enjoys (nearly) the same complexity on querying the operator $F$ as the optimal complexity achieved by projection-based methods, even though the projection oracle is replaced by quadratic programming solvers. For example, when the operator $F(x)$ is monotone, we show that CGM achieves a weak $\epsilon$-solution with $\mathcal{O}(1/\epsilon^{2})$ queries to $F(x)$ and $\mathcal{O}(1/\epsilon^{2})$ calls to a quadratic programming solver. To the best of our knowledge, our algorithm is the first primal method that achieves the optimal complexity on queries to $F$ for functional constrained monotone variational inequality problems without requiring any information on the optimal Lagrange multipliers. This information includes whether the optimal Lagrange multiplier exists and how its magnitude can be estimated. A comparison with existing results can be found in Table 1.
@@
We further illustrate that the quadratic program at each iteration of CGM allows for efficient implementation and permits even closed-form solutions in special cases such as simplex constraints (Algorithm 4) or when there is only one (nonlinear) constraint function (Algorithm 2). This yields a direct implementation of our algorithm comparable to unconstrained methods. Empirically, we evaluate the convergence of CGM through several numerical experiments and demonstrate its effectiveness.
Table 1: Complexity on queries of the operator $F$ to guarantee $\max_{x\in\mathcal{C}}F(x)^{\top}(\hat{x}-x)\leq\epsilon$ (optimality) and $\max_{i}g_{i}(\hat{x})\leq\epsilon$ (feasibility) for the monotone setting, where $\hat{x}$ is the output of the algorithm. ACVI [92, 14] guarantees last-iterate convergence for strong $\epsilon$-solutions $\max_{x\in\mathcal{C}}F(\hat{x})^{\top}(\hat{x}-x)\leq\epsilon$, but it also requires stronger assumptions that either $F$ is strictly monotone or one of $g_{i}$ is strictly convex. In the table, “P” denotes primal methods, while “PD” stands for primal-dual methods. Primal-dual methods require the existence of optimal Lagrange multipliers $\lambda^{*}$, while CGM does not. At each iteration of the algorithm, projected gradient method (PGM) requires access to a projection oracle (PO), Frank-Wolfe method requires access to a linear minimization oracle (LMO), ACVI requires solving a strongly-convex optimization (SCO) sub-problem, and CGM requires solving a simple and possibly sparse quadratic program (QP) $\min_{v\in V_{\alpha}(x_{t})}(1/2)\lVert v+F(x_{t})\rVert^{2}$ (see details in Algorithm 1).
Method Type Optimality Feasibility Per-Iter. Cost Requirement
PGM [23] P $\mathcal{O}(1/\epsilon^{2})$ PO -
Frank-Wolfe [86] P $\mathcal{O}(1/\epsilon^{2})$ LMO -
ACVI [92, 14] PD $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ SCO Existence of $\lambda^{*}$
ConEx [7] PD $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ Existence of $\lambda^{*}$
CGM (Algorithm 1) P $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ QP -
Literature review.   Early advances on iterative methods for solving variational inequality problems can be found in Pang and Chan [73], Facchinei and Pang [23], Noor [72]. Here, we focus on the literature on first-order methods. For monotone variational inequality problems, the projected gradient method achieves the optimal convergence rate $\mathcal{O}(1/\epsilon^{2})$ [23]. If the operator is also Lipschitz, the optimal convergence rate $\mathcal{O}(1/\epsilon)$ has been established for various algorithms, such as the extragradient [66], optimistic gradient [60], and the dual extrapolation method [67]. These results focused on the deterministic setting with guarantees on the average iterates. Recent efforts include extension to the stochastic setting [37, 41, 47, 93, 42, 32, 48, 33, 1] and examination of last-iterate convergence guarantees [26, 12, 11, 29, 30]. Convergence to strong $\epsilon$-solutions was considered in Dang and Lan [17], Diakonikolas [19]. Adaptive and parameter-free algorithms were developed in Malitsky [56], Diakonikolas [19]. There is also a line of research on the relaxation of the monotone assumption, including the quasi-monotone setting [4, 69, 2], the pseudo-monotone setting [83, 42, 9], and the non-monotone setting [84, 91, 50, 20, 74]. All theses works either consider the unconstrained case or assume access to a projection oracle.
For functional constrained problems, previous works [92, 14, 7] centered around primal-dual methods based on the Lagrangian function. Their analysis requires the existence and boundedness of the optimal Lagrange multipliers. Instead, we provide the first primal algorithm, CGM (Algorithm 1), that does not require any information on the optimal Lagrange multipliers. CGM reduces to constrained gradient descent [61] for minimization problems when $F$ is a gradient field. Muehlebach and Jordan [61] only considered smooth and strongly-convex minimization problems, which form a subclass of Lipschitz and strongly-monotone variational inequality problems. Moreover, they use a stepsize dependent on the optimal Lagrange multipliers and require the quadratic program to be exactly solved. In contrast, our analysis of CGM applies to monontone variational inequality problems without the Lipschitz assumption on the operator $F$ and allows certain inaccuracies in the quadratic programming solver. More importantly, no information on the optimal Lagrange multipliers is required.
@@ 2.1 Variational Inequality Problems
In our paper, we assume that $\mathcal{C}$ is contained in an Euclidean ball with diameter $D$. This entails that the convex feasible set $\mathcal{C}$ is also compact. By Corollary 2.2.5 in Facchinei and Pang [23], the solution set of the variational inequality problem (1) is nonempty and compact. We are interested in finding the following weak $\epsilon$-approximate solution of the constrained variational inequality problem (1).
@@
For some $\epsilon>0$, we call a point $\hat{x}\in\mathbb{R}^{d}$ a weak $\epsilon$-solution of the variational inequality problem (1), where the feasible set is of the form (2), if $F(x)^{\top}(\hat{x}-x)\leq\epsilon$, $\forall x\in\mathcal{C}$ and $g_{i}(\hat{x})\leq\epsilon$, $\forall i\in[m]$.
The above definition of the weak $\epsilon$-solution for a variational inequality problem is commonly adopted in the related literature [66, 67, 41, 7]. A strong $\epsilon$-solution, $\hat{x}\in\mathcal{C}$, is characterized by $F(\hat{x})^{\top}(\hat{x}-x)\leq\epsilon$, $\forall x\in\mathcal{C}$. When $\epsilon=0$ and $F$ is monotone and continuous, the two solution sets are equivalent [58, 44]. When $\epsilon>0$ and $F$ is monotone, every strong $\epsilon$-solution is also a weak $\epsilon$-solution.
@@ 3 The Constrained Gradient Method (CGM)
A standard algorithm in the literature on monotone variational inequality problems is the projected gradient method, which requires access to a projection oracle and performs the following updates at each iteration $t=0,1,\ldots,T-1$:
@@
In Algorithm 1, we allow for an inexact solution of the quadratic program, where the stopping criteria can be satisfied using various methods. More details on how this quadratic program can be solved are discussed in Section 5. The auxiliary constraint $g_{m+1}(x)=\lVert x\rVert^{2}-D^{2}$ is used to ensure that both the iterate $x_{t}$ and the projected velocity $v_{t}$ are bounded. It is added specifically to simplify the theoretical analysis and will not be used in actual implementations. The output of Algorithm 1 is the average of iterates, which is a standard strategy in the literature [68, 10].
@@ 4.1 The Monotone Setting
The above theorem implies that for CGM to achieve a weak $\epsilon$-solution, the number of queries to the operator $F(x)$ is at most $\mathcal{O}(1/\epsilon^{2})$, and the number of calls to a quadratic programming solver is at most $\mathcal{O}(1/\epsilon^{2})$. Notably, the complexity of querying the operator $F$ matches that of projection-based methods [68] or Frank-Wolfe type methods [86], known to be optimal even for simple constraint sets according to the lower complexity bound (Theorem 3.2.1 in Nesterov [68]).
In stark contrast, the convergence analysis of previous primal-dual methods relies on the information about the optimal Lagrange multipliers. For instance, the ADMM-based interior point method [92, 14] achieves strong $\epsilon$-solutions with the same $\mathcal{O}(1/\epsilon^{2})$ queries to $F$ and requires $\mathcal{O}(1/\epsilon^{2})$ calls to a program that solves a sequence of strongly-convex minimization problems. However, their results require the strong assumption that either $F$ is strictly-monotone or one of the $g_{i}(x)$ is strictly-convex. The constraint extrapolation (ConEx) method [7, 8] achieves a weak $\epsilon$-solution as in Definition 1 with $\mathcal{O}(1/\epsilon^{2})$ calls to the operator $F$. However, they also require the existence and boundedness of the optimal Lagrange multipliers. To the best of our knowledge, CGM is the first primal method that achieves the optimal complexity on queries to $F$ for functional constrained monotone variational inequality problems without any information on the optimal Lagrange multipliers.
@@ 4.2 The Strongly-Monotone Setting
When $F(x)$ is $\mu$–strongly-monotone, i.e., $(F(x_{t})-F(x))^{\top}(x_{t}-x)\geq\mu\lVert x_{t}-x\rVert^{2}$, we obtain that $\forall\,x\in\mathcal{C}$,
@@
$\begin{split}(t+2)^{c_{\gamma}}g_{i}(x_{t+1})&\leq\left(1-\frac{c_{\gamma}}{t+1}\right)(t+2)^{c_{\gamma}}g_{i}(x_{t})+\frac{c_{M}(t+2)^{c_{\gamma}}}{(t+1)^{2}}\\ &=\left(1-\frac{c_{\gamma}}{t+1}\right)\left(1+\frac{1}{t+1}\right)^{c_{\gamma}}(t+1)^{c_{\gamma}}g_{i}(x_{t})+\frac{c_{M}}{(t+1)^{2-c_{\gamma}}}\left(1+\frac{1}{t+1}\right)^{c_{\gamma}}\\ &\leq\left(1-\frac{c_{\gamma}}{t+1}\right)\left(1+\frac{c_{\gamma}}{t+1}\right)(t+1)^{c_{\gamma}}g_{i}(x_{t})+\frac{c_{M}}{(t+1)^{2-c_{\gamma}}}\left(1+\frac{c_{\gamma}}{t+1}\right)\\ &\leq(t+1)^{c_{\gamma}}g_{i}(x_{t})+\frac{2\,c_{M}}{(t+1)^{2-c_{\gamma}}},\end{split}$ (9)
@@
$(\kappa(t+1)+2)^{c_{\gamma}}g_{i}(x_{\kappa(t+1)+1})$ $\leq(\kappa(t+1)+2)^{c_{\gamma}}\max\left\{\frac{ML_{g}}{\kappa(t+1)+1},\frac{c_{M}}{(\kappa(t+1)+1)^{2}}\right\}$
$=\left(1+\frac{1}{\kappa(t+1)+1}\right)^{c_{\gamma}}\max\left\{\frac{ML_{g}}{(\kappa(t+1)+1)^{1-c_{\gamma}}},\frac{c_{M}}{(\kappa(t+1)+1)^{2-c_{\gamma}}}\right\}$
$\leq 2\max\left\{\frac{ML_{g}}{(\kappa(t+1)+1)^{1-c_{\gamma}}},\frac{c_{M}}{(\kappa(t+1)+1)^{2-c_{\gamma}}}\right\}$
$\leq 2\max\{ML_{g},c_{M}\}.$
@@
In the strongly-monotone setting, the convergence rate of CGM in terms of the optimality gap is $\mathcal{O}(1/T)$, which matches the rate of projection-based methods [68]. However, it is worth noting that the rate in terms of the constraint violation approaches $\mathcal{O}(1/T)$ only as $\gamma$ grows large enough. We leave the task of closing this gap to future work. As a comparison, the primal-dual method in Yang et al. [92] achieves $\mathcal{O}(1/\sqrt{T})$ rate measured by the distance $\lVert\bar{x}_{T}-x^{*}\rVert$. If the operator is additionally Lipschitz, the same $\mathcal{O}(1/\sqrt{T})$ rate is attained for a strong approximate solution.